OpenStudy (anonymous):

Which number is between 0.5 and 8/9???

OpenStudy (freckles):

\[0.5=\frac{1}{2}=\frac{1}{2} \cdot \frac{\frac{9}{2}}{\frac{9}{2}}=\frac{9/2}{9}=\frac{4.5}{9}\] so you are asking what number is between 4.5/9 and 8/9 well if the denominator is 9 then all you have to do is compare the numerators is the numerator between 4.5 and 8?

OpenStudy (anonymous):

is it 22/24?? @freckles

OpenStudy (anonymous):

it's suppose to be a fraction

OpenStudy (anonymous):

@freckles ya there?

OpenStudy (anonymous):

oh its 25/36

OpenStudy (freckles):

ok it doesn't look like any of you choices are over 9

OpenStudy (anonymous):

0.5 = 18/36 8/9 = 32/36 18/36 + 32/36 = 50/36 50/36 ÷ 2 = 25/36 25/36 is the number in between 0.5 and 8/9

OpenStudy (freckles):

that sounds like a lovely answer actually

OpenStudy (anonymous):

thankyou can you help me with 4 more please c;

OpenStudy (freckles):

so your question was just which of the choices is between 0.5 and 8/9 right? (I'm kinda curious as to the other operations going on above)

OpenStudy (anonymous):

oh okay XD thats okay i dont need help c:

OpenStudy (freckles):

you don't need help anymore?

OpenStudy (anonymous):

oh wait actually i need help with a few more if thast okay.

OpenStudy (freckles):

ok

OpenStudy (anonymous):

Express 3/13 as a decimal.

OpenStudy (freckles):

to what digit or we to round?

OpenStudy (anonymous):

it doesn't say

OpenStudy (freckles):

 numerator (top number) goes inside the box

OpenStudy (freckles):

you can write 3 as 3.how ever many zeros you need or what 3.000000

OpenStudy (freckles):